# Unsourced Random Access: A Comprehensive Survey

### v2

@@ VI-D Integrated Sensing and Communication in URA
In [154], the problem of integrating sensing and communication is considered in a URA setup, where a massive number of unsourced and sporadically active users are connected to the system. In the resulting system, called unsourced ISAC (UNISAC), users are divided into two main categories: Sensing and communication users. The system aims to decode signals from communication users while detecting signals from active sensing users and estimating their AOAs. The authors derive an achievable bound for the UNISAC to show its efficacy in sensing the sensing users as well as in communicating with the communication ones in a URA setup.
There are still many aspects of an ISAC system in a URA setup that need further investigation. For instance, the problem becomes more challenging when sensing users are passive, as opposed to the active assumption in [154]. In this case, only communication users transmit signals to the BS, with the same signal opportunistically used to sense passive targets.
@@
URA-specific challenges: In a URA system, a massive number of users transmit sporadically without being registered or authenticated by the network. Therefore, there is no control over transmitter characteristics such as location, signal structure, and beam direction, which are critical for reliable passive target sensing. Additionally, the random access nature of user transmissions causes heavy interference, a phenomenon uncommon in traditional sensing techniques, resulting in a lack of solutions to mitigate it. Therefore, advanced algorithms are needed to address these challenges and enhance the system’s ability to sense passive targets effectively.
VI-E URA with Feedback
@@
In communication systems, feedback has been shown to greatly simplify transmission and reception mechanisms and, in some cases, even improve system performance [155]. The main challenge in employing feedback-based techniques in URA is the lack of coordination between the users and the receiver due to the massive number of potential users. However, even in URA, there can be some level of coordination as long as the signal sent from the receiver is not user-specific, as there is no user identity in URA. For instance, the receiver can broadcast a common pilot/beacon to the users so that they check if they are in the set of successfully decoded users in the previous round, and estimate their channel coefficients or at least their large-scale components depending on their dedicated time slots. Just note that the time dedicated to this operation needs to be considered as extra overhead when comparing with the solutions without feedback. This issue is partly tackled in [156, 157, 158]. In [156], hashes of the messages of successfully decoded users and their estimated channel coefficients are fed back using beamforming techniques; while in [157], with the help of a designed feedback beacon, the users estimate their channels and infer whether they were successfully decoded using a pre-defined threshold. One application of URA can be in sensor networks where the devices partially observe a common state, which is the main interest of the receiver. For such an application, the authors in [158] devise a mechanism where the receiver broadcasts a reliability indicator of the estimated state after each transmission, using which each device can decide whether to transmit in the next iteration.
@@
Recall that the main novelty of URA is providing a unified framework to decouple user identification and data transmission problems. This makes it a fit solution in applications where user identification is not feasible at all due to the massive number of potential users. However, URA is still applicable and a good candidate in non-massive (but with still a considerable number of potential users) scenarios where user identification is still feasible. In these scenarios, conventional authentication and acknowledgment schemes impose a huge overhead on the system. This issue is tackled in [159].
VI-F RIS-Aided URA
@@
In URA, due to the massive number of potential users, it is very probable that the line-of-sight link between some users and the receiver is either not present or heavily attenuated. In this case, employing relaying techniques such as RIS can play an important role in improving the overall system performance by providing a substitute communication link with potentially many degrees of freedom. However, the huge channel estimation overhead is the main challenge in the use of RIS in URA systems. The channel estimation overhead is already one of the main issues in RIS systems due to the high number of channel coefficients per user as a result of the high number of RIS elements and the potentially massive number of antennas at the receiver (as the benefits of a RIS system are amplified when combined with a massive MIMO receiver). This problem is more pronounced in a RIS-aided URA system as there is a potentially massive number of even users, which renders most existing channel estimation schemes in RIS infeasible. There are a few works that have considered the RIS-aided URA scenario [123, 124, 125]. Although these works make the channel estimation feasible in a RIS-aided URA setup by employing an SIC-based iterative approach and time frame slotting strategies, the high computational complexity still remains a challenge.
@@
Moreover, there are many promising RIS configurations and types that need to be investigated in a URA setup, such as active or hybrid RIS [160] and metasurface-based RIS [161], in order to find suitable candidates. Moreover, the employment of massive arrays such as extra large massive MIMO and RIS leads to some active users being in the near field of these arrays, which makes hybrid near field-far field analysis a necessity. This is in contrast with the existing URA solutions, which are all in the far-field regime.
VI-G Standardization
@@
There has not been any effort in the standardization of URA, neither for incorporating it in the next updates of the current protocol nor for future generations such as 6G. This requires a thorough investigation and analysis of the current competing solutions, and finding a good candidate that preferably can support both the current protocols, perhaps with small modifications, and the high demands of the future ones. For instance, in [162], with a small modification in the 5G new radio (5G-NR) release 16 protocol, the authors achieve an improvement in energy efficiency of about 5 to 10 dB compared to the existing 5G-NR solution for systems with a small number of active users (10 to 50 users), employing an IDMA-based URA solution. However, this solution is not scalable to higher active user loads.
@@
URA-specific challenges: There are two main challenges in incorporating URA in current protocols, such as 5G-NR and beyond, for reasonable active user loads (e.g., in the order of hundreds). The first is the small number of preamble sequences available in 5G-NR as opposed to the massive number of preamble sequences (pilots) needed in the preamble/pilot-based URA solutions. The second is the limited number of access patterns that can be realized, which is vital in transmission pattern-based URA solutions such as ODMA-based ones. This shows that an important step in the realization of URA in future protocols is to considerably increase the available pilot or transmission patterns.

### v3

@@ VI-B5 ISAC in URA
The problem of ISAC in URA is introduced for the first time in [171], where a massive number of unsourced and sporadically active users are connected to the system to jointly perform sensing and communication tasks. In the resulting system, called unsourced ISAC (UNISAC), users are divided into two main categories: Sensing and communication users. The system aims to decode signals from communication users while detecting signals from active sensing users and estimating their AOAs. The authors derive an achievability bound for UNISAC to demonstrate its effectiveness in enabling both user sensing and user communication within a URA setup. This work is conducted under two key assumptions: the targets are active, and the channels between the users (targets) and the BS are line-of-sight (LoS).
The problem of UNISAC has been further investigated in subsequent works, including: proposing a practical transceiver scheme for UNISAC under LoS channels between the BS and all sensing and communication users [173, 174], as well as under LoS target–BS channels and Rayleigh fading user–BS channels [175]; studying the UNISAC model equipped with fluid antenna systems [176, 177]; and addressing the problem of passive target sensing [180].
There are several promising directions for future research. In existing UNISAC setups, only the problem of user angle estimation has been investigated, whereas the field of sensing is much broader. For example, tracking, surveillance, passive source localization, and target localization are important sensing tasks that warrant further study. Moreover, networks incorporating advanced technologies, such as multi-BS systems, UAV-equipped systems, and RIS-aided systems, can be explored to enhance performance. Another significant research direction is to study UNISAC under realistic channel models, including frequency-selective fading, large-scale fading, Doppler effects, asynchronous transmission errors, and hardware impairments [172].
@@
To explore the challenges that future work on URA-based ISAC may encounter, many aspects still require further investigation. In a URA system, a massive number of users transmit sporadically without being registered or authenticated by the network. Therefore, there is no control over transmitter characteristics such as location, signal structure, and beam direction, which are critical when the transmitters’ echoes are used for reliable passive target sensing. Additionally, the random access nature of user transmissions causes heavy interference, a phenomenon uncommon in traditional sensing techniques. Advanced algorithms are necessary to address these challenges and enhance the system’s ability to effectively sense passive targets.
VI-B6 URA with Feedback
@@
In communication systems, feedback has been shown to greatly simplify transmission and reception mechanisms and, in some cases, even improve system performance [178]. The main challenge in employing feedback-based techniques in URA is the lack of coordination between the users and the receiver due to the massive number of potential users. However, even in URA, there can be some level of coordination as long as the signal sent from the receiver is not user-specific, as there is no user identity in URA. For instance, the receiver can broadcast a common pilot or beacon to the users, allowing them to check if they are among the successfully decoded users from the previous round and estimate their channel coefficients or at least their large-scale components, depending on their dedicated time slots. Note that the time dedicated to this operation needs to be considered as extra overhead when comparing with solutions without feedback. This issue is tackled in part in [180, 179, 181, 182]. In [179], hashes of the messages of successfully decoded users and their estimated channel coefficients are fed back using beamforming techniques. In [180], the authors enhance the approach of [179] by informing both the successfully decoded users and those that were not decoded about their respective transmission outcomes. In [181], with the help of a designed feedback beacon, the users estimate their channels and infer whether they were successfully decoded using a predefined threshold. An application of URA can be in sensor networks where the devices partially observe a common state, which is the main interest of the receiver. For such an application, the authors in [182] devise a mechanism where the receiver broadcasts a reliability indicator of the estimated state after each transmission, using which each device can decide whether to transmit in the next iteration.
@@
Recall that the main novelty of URA is providing a unified framework to decouple user identification and data transmission problems. This makes it a suitable approach in applications where user identification is not feasible due to the massive number of potential users. However, URA remains applicable and is a good candidate in non-massive scenarios with a considerable number of potential users, where user identification is still feasible. In these scenarios, conventional authentication and acknowledgment schemes impose a huge overhead on the system. This issue is tackled in [183].
VI-B7 RIS-Aided URA
@@
In URA, due to the massive number of potential users, it is very probable that the line-of-sight link between some users and the receiver is either not present or heavily attenuated. In this case, employing relaying techniques such as RIS can play a crucial role in enhancing the overall system performance by providing an alternative communication link with potentially many degrees of freedom. However, the huge channel estimation overhead is the main challenge in the use of RIS in URA systems. The channel estimation overhead is already one of the main issues in RIS systems due to the high number of channel coefficients per user as a result of the large number of RIS elements and the potentially massive number of antennas at the receiver (as the benefits of a RIS system are amplified when combined with a massive MIMO receiver). This problem is more pronounced in a RIS-aided URA system as there is a potentially massive number of even users, which renders most existing channel estimation schemes in RIS infeasible. There are a few works that have considered the RIS-aided URA scenario [136, 137, 138]. Although these works make channel estimation feasible in a RIS-aided URA setup employing an iterative SIC-based approach and time frame slotting strategies, the high computational complexity remains a significant challenge.
@@
Moreover, there are many promising RIS configurations and types that need to be investigated in a URA setup, such as active or hybrid RIS [184] and metasurface-based RIS [185], in order to find suitable candidates. Moreover, the employment of massive arrays such as extra-large massive MIMO and RIS leads to some active users being in the near-field of these arrays, which makes hybrid near-field-far-field analysis necessary. This is in contrast with the existing URA solutions, which are all for the far-field regime.
VI-B8 Security
One of the key concerns in multiple access systems is a robust implementation against security threats, such as jamming and eavesdropping. To address these security threats, some common countermeasures are employed, including mutual authentication protocols for access control and authentication, such as the authentication and key agreement (AKA) used in 4G/5G, and data encryption techniques, such as AES-128 in 4G and AEGIS-256 in 5G. These protocols ensure that both users and the receiver (network) verify each other’s identities before establishing a secure connection. However, many of these solutions are not feasible in a URA setup, primarily due to the lack of coordination between users and the receiver. For example, key agreement in the AKA protocol relies on some level of coordination between the user and the receiver to establish a secure connection.
To the best of our knowledge, there has not been any work addressing the security aspects of URA. However, some existing security solutions can still be utilized in URA with certain modifications or considerations. For example, light encryption methods can be employed (similarly to those in [186] and [187]) to introduce a certain level of security with a common key among users. In some URA solutions, the key can also be randomly selected from a pool, with each key paired with a transmission signature, such as a pilot sequence in pilot-based schemes and a transmission pattern in ODMA-based ones. Likewise, physical layer security techniques, such as the injection of artificial noise, can also be employed, with the modification that all users utilize the same noise sequence or randomly select it from a pool of available sequences. Existing physical layer authentication techniques based on signal features, such as channel state information or RF fingerprints [188], can also be utilized as long as they do not rely on user-specific signatures.
VI-B9 Standardization
@@
There has been no effort in standardizing URA so far. This requires a thorough investigation and analysis of the current competing solutions, and finding a good candidate that can preferably support both the current protocols, perhaps with minor modifications, and the high demands of the future ones. For instance, in [189], with a slight modification in the 5G new radio (5G-NR) release 16 protocol, the authors achieve an improvement in energy efficiency of about 5 to 10 dB compared to the existing 5G-NR solution for systems with a small number of active users (10 to 50 users), employing an IDMA-based URA solution. However, this solution is not scalable to higher active user loads.
@@
There are two main challenges in incorporating URA in current protocols, such as 5G-NR and beyond, for reasonable active user loads (e.g., in the order of hundreds). The first is the small number of preamble sequences available in 5G-NR as opposed to the massive number of preamble sequences (pilots) needed in the preamble/pilot-based URA solutions. The second is the limited number of access patterns that can be realized, which is vital in transmission pattern-based URA solutions such as ODMA-based ones. This indicates that a crucial step in realizing URA in future protocols is to considerably increase the available pilot or transmission patterns.
This comprehensive survey of URA reveals an active field that has evolved significantly in recent years, with substantial progress in fading and MIMO channels. The research challenges identified above demonstrate both the complexity of remaining questions and the rich opportunities for future investigation. Addressing these challenges will be essential for the practical adoption and deployment of URA solutions in future wireless systems.